OpenStudy (thats_kyy):

Simplify the following expression: 2x − 4y + 3x2 + 7y − 16x. 3x2 − 14x − 3y 3x2 + 14x + 3y 3x2 − 18x + 3y 3x2 − 14x + 3y

OpenStudy (thats_kyy):

@nn

OpenStudy (thats_kyy):

@Nnesha

OpenStudy (johnn.):

Let's simplify step-by-step. 2x−4y+3x2+7y−16x =2x+−4y+3x2+7y+−16x Combine Like Terms: =2x+−4y+3x2+7y+−16x =(3x2)+(2x+−16x)+(−4y+7y) =3x2+−14x+3y Answer: =3x2−14x+3y